OpenStudy (idealist10):

Find the general solution of x^2*y"+xy'-y=4/x^2, given that y1=x satisfies the complementary equation.

OpenStudy (idealist10):

\[x ^{2}y''+xy'-y=\frac{ 4 }{ x^2 }\]

OpenStudy (idealist10):

y=ux y'=u'x+u y"=u"x+2u' ------------ Subbing: \[x^2y''+xy'-y=x^2(u''x+2u')+x(u'x+u)-ux=\frac{ 4 }{ x^2 }\]

OpenStudy (idealist10):

Simplify, I got \[x^3u''+3x^2u'=\frac{ 4 }{ x^2 }\]

OpenStudy (idealist10):

Substitution: z=u' x^3*z'+3x^2*z=4/x^2 z'+(3/x)z=4/x^5 integrating factor is x^3 so d/dx(x^3*z)=4/x^2 x^3*z=integral of 4/x^2 dx x^3*z=-4/x+C so \[z=-\frac{ 4 }{ x^4 }+\frac{ C }{ x^3 }\]

OpenStudy (idealist10):

Since z=u', u'=-4/x^4+C/x^3 integrate u', \[u=\frac{ 4 }{ 3x^3 }-\frac{ C _{2} }{ 2x^2 }+C _{1}\]

OpenStudy (idealist10):

\[y=\frac{ 4 }{ 3x^2 }+C _{1}x+\frac{ C _{2} }{ 2x }\]

OpenStudy (idealist10):

But the book's answer is y=4/3x^2+C1*x+C2/x. So I don't know where I got wrong. Please tell me where my mistake is.

OpenStudy (idealist10):

@Directrix @pooja195 @surjithayer @Nnesha @Preetha

OpenStudy (idealist10):

@Kainui

OpenStudy (kainui):

They look the same to me, since \(\frac{C_2}{2}\) is just as arbitrary of a constant as plain \(C_2\) they just got rid of the 2 because it doesn't change the answer.

OpenStudy (idealist10):

So you're saying that \[\frac{ C _{2} }{ 2x }=\frac{ C _{2} }{ x }\]?

OpenStudy (idealist10):

And integrating C/x^3 gives you -C/2x^2 but it's also the same thing as C/2x^2, right? Since you said the term got absorbed into the constant C, right? So the book's answer is correct?

OpenStudy (kainui):

Yeah, the book is right and you are right. One thing though, actually using the same letter on both sides makes it false, because if you were to solve that \[\frac{C_2}{2x} = \frac{C_2}{x}\] you'd get that: \[C_2 = 2*C_2\] which implies \(C_2=0\) which is not what you mean. So it's technically correct to write something to distinguish them apart, like use a different letter. \[\frac{C_2}{2x} = \frac{D_2}{x}\] This is fine, cause all this says is \(C_2 = 2* D_2\). But no arbitrary constant is better than any other arbitrary constant so w/e

OpenStudy (idealist10):

Okay, thank you so much for pointing out. :)